# Segment Anything Model (SAM) for Radiation Oncology

### v1

@@ 3.3 Automatic segmentation of everything
In the ’segment everything’, SAM is designed to create segmentation masks for every possible object present within the full image, no manual priors are needed. This mode is considered the first testing approach. The commencement of this method involves producing a grid of point prompts, also known as grid sampling, which spans the entire image. Following that, the prompt encoder uses the uniformly sampled grid points to generate point embeddings, which are then merged with the image embeddings. The mask decoder then receives this blend as input and delivers multiple potential masks for the entire image. Afterward, a filter system is put into action to eliminate duplicate and inferior masks. For the segment everything mode, we name it SE for corresponding results analysis.
@@ 3.4 Manual box prompt
In prompt mode, the box prompt signifies the spatial region that necessitates segmentation, representing the object of interest. This principal mode of evaluation employs prompts that mimic a human user’s interaction, crafted while closely observing the objects. Our focus lies primarily on the box prompt, tailored to encapsulate SAM’s realistic use cases for creating image masks. An experienced medical physicist typically places the box prompt, guided by anatomical characteristics and clinical experience. The placement is usually in close proximity to the Region of Interest (ROI) margin. It’s vital to remember that a single "object" of interest or a "ground truth" mask might comprise multiple disconnected segments, a situation commonly encountered in medical imaging. To ensure each distinct, contiguous region of the object of interest is accurately represented, multi-box prompts are strategically placed.
@@ 3.5 Evaluation metrics
In order to thoroughly assess SAM’s segmentation performance, we employed two commonly used metrics, as detailed below: 1) Dice Coefficient (DICE, %): This measure of similarity is used to evaluate the degree of overlap between the prediction and the ground truth (GT). With a range between [0, 1], a higher value denotes a more successful performance by the model.
2) Jaccard Similarity Coefficient (JAC, %): Also recognized as the Intersection over Union (IOU), this metric, although similar to DICE, poses more stringent demands. It quantifies the complete overlap of label ensembles across multiple test images, accommodating fractional labels through the application of fuzzy set theory. Like the DICE coefficient, higher JAC values signify superior model performance.
Figure 1: Comparison of the clinical delineation, SAM segment everything, and SAM with manual box prompt of example cases. a) prostate case b) lung case c) gastrointestinal d) head&neck
@@ 4.1 Example cases
Figure 1. shows the auto-segmented contours from the two experiments (SAM segment everything and SAM box prompt) and the clinical delineation in the axial plane of one typical prostate (Fig. 1(a)) case, one typical lung (Fig. 1(b)) case, one typical gastrointestinal (Fig. 1(c)) case, and one typical head & neck (Fig. 1(d)) case.
@@ 4.2 SAM model accuracy
For the prostate region, as shown in Figure 2 and Figure 4, the SAM (Segment Everything Mode) resulted in a Dice score of 0.3±0.06 and a JAC (Jaccard Index) score of 0.18±0.04 when outlining the prostate. In the bladder adjacent to the prostate, the Dice score was 0.49±0.05 and the JAC score was 0.33±0.04. For the femoral head_L and femoral head_R, their Dice scores were both around 0.4, and JAC scores were around 0.25. The Dice score for the rectum was relatively low, at 0.19±0.07, and the JAC score was 0.11±0.03.
For the lung region, as shown in Figure 2 and Figure 4, SAM (Segment Everything Mode) produced Dice scores around 0.39 for both lung_L and lung_R, and JAC scores around 0.26. The spinal cord had a Dice score of 0.21±0.05, and a JAC score of 0.11±0.02. The heart had a relatively low Dice score of 0.14±0.03, and a JAC score of 0.07±0.01. For the esophagus, SAM Segment Everything mode was unable to outline or recognize it, resulting in Dice and JAC scores of around 0.
For the GI (gastrointestinal) region, as shown in Figure 3 and Figure 5, SAM Segment Everything Mode produced a Dice score of 0.41±0.06 and a JAC score of 0.26±0.05 when outlining the liver. The results for kidney_L and kidney_R were similar, with Dice scores around 0.5 and JAC scores around 0.33. The spinal cord results were consistent with those of the lung region, with Dice and JAC scores of about 0.2 and 0.1 respectively. The Dice score for the stomach was relatively low, at 0.1±0.02, and the JAC score was 0.05±0.01. For the large bowel and small bowel, SAM Segment Everything mode was unable to outline or recognize them, resulting in Dice and JAC scores of around 0.
For the head & neck region, as shown in Figure 3 and Figure 5, SAM Segment Everything Mode produced a Dice score of 0.49±0.06 and a JAC score of 0.3±0.05 when outlining the brain. The mandible had a Dice score of 0.38±0.09, and a JAC score of 0.23±0.07. The spinal cord had similar results to other regions, with Dice and JAC scores of about 0.2 and 0.1 respectively. For the parotid_L, parotid_R, cochlea_L, and cochlea_R, SAM Segment Everything mode was unable to outline or recognize them, resulting in Dice and JAC scores of around 0.
@@ 4.3 Impact of box prompt
Overall, after introducing the manual box prompt, there was an improvement in Dice and JAC results for most organs. Some organs that could not be identified in SAM Segment Everything Mode were now recognizable, although the resulting Dice and JAC values were low. However, a few small organs in the head and neck area remained unrecognizable even after the box prompt was added.
@@
For the prostate region, as shown in Figure 2 and Figure 4, the manual box prompt mode resulted in a Dice score of 0.41±0.05 and a JAC score of 0.25±0.03 when outlining the prostate. The bladder had a Dice score of 0.59±0.07, and a JAC score of 0.43±0.06. For femoral head_L and femoral head_R, their Dice scores each increased by around 0.1, and their JAC scores each increased by about 0.09. The rectum’s Dice score increased to 0.29±0.05, and the JAC score to 0.17±0.03.
@@
For the lungs, as depicted in Figure 2 and Figure 4, manual box prompt mode improved the Dice scores for both lung_L and lung_R by about 0.11, and the JAC scores by about 0.07. The spinal cord had a Dice score of 0.5±0.09, and a JAC score of 0.33±0.04. The heart had a relatively modest improvement, with Dice scores and JAC scores increasing by about 0.05 and 0.03, respectively. The esophagus was recognizable under manual box prompt mode, but both Dice and JAC scores were low, at about 0.14 and 0.07, respectively.
@@
For the GI region, as shown in Figure 3 and Figure 5, the manual box prompt mode yielded a Dice score of 0.49±0.04 and a JAC score of 0.32±0.03 when outlining the liver. For kidney_L and kidney_R, the Dice and JAC scores both increased by about 0.1 and 0.08 respectively. The spinal cord saw an increase of around 0.3 for Dice score and 0.22 for the JAC score. For the stomach, the Dice and JAC scores increased by about 0.07 and 0.05 respectively. The large bowel was recognizable under manual box prompt mode, but both Dice and JAC scores were low, at around 0.1 and 0.05 respectively. For the small bowel, manual box prompt mode was still unable to recognize it, with both Dice and JAC values being around 0.
@@
For the head and neck region, as shown in Figure 3 and Figure 5, the manual box prompt mode resulted in a Dice score of 0.59±0.06 and a JAC score of 0.43±0.07 when outlining the brain. The mandible had a Dice score of 0.51±0.11 and a JAC score of 0.3±0.04. The spinal cord had similar improvements as other regions, with increases of around 0.29 and 0.23 for the Dice and JAC scores respectively. For parotid_L and parotid_R, they were recognizable under manual box prompt mode, but both Dice and JAC scores were low, at around 0.1 and 0.09 respectively. For cochlea_L and cochlea_R, manual box prompt mode was still unable to recognize them, with both Dice and JAC values being around 0.
@@
Figure 2: Boxplot (minimum, first quartile, median, third quartile, and maximum, respectively) of Dice coefficients of ROIs between the ground clinical delineation and the SAM auto-segmented from three different experiments of the testing cases for prostate and lung. SAM segment everything and SAM manual box prompt correspond to 1 (grey) and 2 (red) in the figure.
@@
Figure 3: Boxplot (minimum, first quartile, median, third quartile, and maximum, respectively) of Dice coefficients of ROIs between the ground clinical delineation and the SAM auto-segmented from three different experiments of the testing cases for gastrointestinal and head&neck. SAM segment everything and SAM manual box prompt correspond to 1 (grey) and 2 (red) in the figure.
@@
Figure 4: Boxplot (minimum, first quartile, median, third quartile, and maximum, respectively) of Jaccard coefficients of ROIs between the ground clinical delineation and the SAM auto-segmented from three different experiments of the testing cases for prostate and lung. SAM segment everything and SAM manual box prompt correspond to 1 (grey) and 2 (red) in the figure.
@@
Figure 5: Boxplot (minimum, first quartile, median, third quartile, and maximum, respectively) of Jaccard coefficients of ROIs between the ground clinical delineation and the SAM auto-segmented from three different experiments of the testing cases for gastrointestinal and head&neck. SAM segment everything and SAM manual box prompt correspond to 1 (grey) and 2 (red) in the figure.
@@ 5 Discussion
In this section, we analyze the current performance of segmentation in radiation oncology and expand on the possible future perspectives, practical applications, and prospective improvements of SAM in the realm of radiation oncology, based on the detailed results and insights from our study.
@@ 5.1 Current performance in radiation oncology-related segmentation
As for the Dice and Jaccard results, under the SAM "segment everything" mode, the autosegmentation outcomes were satisfactory for the prostate region’s prostate, bladder, femoral head_L, and femoral head_R, while they were less desirable for the rectum. For the lung region, autosegmentation for the lung_L, lung_R, and spinal cord was relatively better, yet less favorable for the heart, with the esophagus not being recognized for segmentation. The auto-segmentation outcomes for the gastrointestinal region’s liver, kidney_L, kidney_R, and spinal cord were relatively better, whereas those for the stomach were less satisfactory, with the large bowel and small bowel not recognized for segmentation. For the head&neck region, the brain, mandible, and spinal cord were better segmented, whereas the parotid_L, parotid_R, cochlea_L, and cochlea_R were not recognized for segmentation. When comparing the results of segmentation across different regions, the "segment everything" mode in SAM performs better for the prostate and lung regions, but less satisfactory for the gastrointestinal and head & neck regions. If the volume and clarity of an organ are taken into account, it can be observed that the model performs better at delineating organs with clear boundaries and larger volumes, such as the liver and brain, and less satisfactory for organs with indistinct boundaries and smaller volumes, such as the parotid and cochlea, which is in general agreement with the experiences of manual clinical delineation. The aforementioned results demonstrate that SAM’s performance in radiotherapy auto-delineation mirrors the clinical experience of human delineation, considering the variation across different regions and organs. Importantly, SAM’s auto-delineation is achieved by a well-trained model, suggesting that SAM exhibits robust generalizability for automatic delineation in radiotherapy.
@@
Upon the inclusion of the box prompt, SAM’s performance in auto-delineation for radiotherapy showed further improvement. For most organs within the four reference regions, Dice values have risen by 0.1 to 0.2, and previously unrecognized organs such as the esophagus and parotid could be identified. However, some organs with vague boundaries, such as the large bowel, small bowel, and cochlea, were still not recognized. This suggests that the box prompt, overall, is effective in enhancing SAM’s performance in radiotherapy segmentation, and future research could consider employing diverse prompt methods for further improvement. SAM is sensitive to the clarity of the Organ at Risk (OAR) boundaries in radiotherapy delineation images; for some organs with unclear boundaries, multi-modality images such as MRI could be considered to assist in the segmentation of some OARs.
@@
Given the existing results, the SAM model demonstrates generalization capabilities consistent with manual radiotherapy delineation, with a single model capable of executing all included test delineations for radiotherapy, and providing an effective prompt technology path for enhancing the performance of radiotherapy delineation. However, judging from the Dice results, SAM has not yet met the accuracy required for clinical radiotherapy. It is generally believed that to achieve acceptable delineation accuracy, the Dice value should reach at least 0.7 when compared with manual clinical delineation [76, 77]. Yet considering that the training of the SAM model largely relies on daily computer vision data, rather than professional clinical imaging data, the results are still acceptable, if not quite impressive [16]. This also suggests that future work can further fine-tune the SAM model in radiotherapy auto-segmentation through the incorporation of more clinical imaging data.

### v2

@@ 3.2 Automatic segmentation of everything
In the ’segment everything’ mode, SAM is designed to create segmentation masks for every possible object present within the full image, no manual priors are needed. This mode is considered the first testing approach. The commencement of this method involves producing a grid of point prompts, also known as grid sampling, which spans the entire image. To enhance the segmentation of the target regions, more random point prompts will be assigned to the target regions guiding an improved segmentation process in this study. Following that, the prompt encoder uses the sampled grid points to generate point embeddings, which are then merged with the image embeddings. The mask decoder then receives this blend as input and delivers multiple potential masks for the entire image. Afterward, a filter system is put into action to eliminate duplicate and inferior masks.
@@ 3.3 Manual box prompt
In the prompt mode, the box prompt signifies the spatial region that necessitates segmentation, representing the object of interest. This principal mode of evaluation employs prompts that mimic a human user’s interaction, crafted while closely observing the objects. Our focus lies primarily on the box prompt, tailored to encapsulate SAM’s realistic use cases for creating image masks. An experienced medical physicist typically places the box prompt, guided by anatomical characteristics and clinical experience. The placement is usually in close proximity to the region of interest margin. It’s vital to remember that a single "object" of interest or a "ground truth" mask might comprise of multiple disconnected segments, a situation commonly encountered in medical imaging. To ensure each distinct, contiguous region of the object of interest is accurately represented, multi-box prompts are strategically placed.
@@ 3.4 Evaluation metrics
In order to thoroughly assess SAM’s segmentation performance, we employed two commonly used metrics, as detailed below:
1) Dice Coefficient (DICE, %): This measure of similarity is used to evaluate the degree of overlap between the prediction and the ground truth (GT) as defined as Eq. 1. With a range between [0, 1], a higher value denotes a more successful performance by the model.
$DICE\left(y,\widetilde{y}\right){=\frac{2\left|y\cap\widetilde{y}\right|}{\left|y\right|+\left|\widetilde{y}\right|}}$ (1)
2) Jaccard Index (JAC, %): Also recognized as the Intersection over Union (IOU), this metric, although similar to DICE, poses more stringent demands as defined as Eq. 2. It quantifies the complete overlap of label ensembles across multiple test images, accommodating fractional labels through the application of fuzzy set theory. Like the DICE coefficient, higher JAC values signify superior model performance.
$JAC\left(y,\widetilde{y}\right)=\frac{\left|y\cap\widetilde{y}\right|}{\left|y\cup\widetilde{y}\right|}$ (2)
where y denotes the volume of clinical manual delineation, and $\widetilde{y}$ denotes the volume of SAM auto-segmentation.
Figure 1: Workflow of SAM segmentation in clinical cases with SAM segment everything mode and SAM box prompt mode taking prostate as an example.
@@ 4.1 Example cases
Figure 2 shows the auto-segmented contours from the two experiments (SAM segment everything and SAM box prompt) and the clinical delineation in the axial plane of one typical prostate (Fig. 2(a)) case, one typical lung (Fig. 2(b)) case, one typical gastrointestinal (Fig. 2(c)) case, and one typical head & neck (Fig. 2(d)) case.
Figure 2: Comparison of the clinical delineation (ground truth), SAM segment everything, and SAM with manual box prompt of example cases. a) prostate b) lung c) gastrointestinal d) head&neck
@@ 4.2 SAM model accuracy
For the prostate, as shown in Figure 3 and Figure 5, the SAM segment everything mode resulted in a Dice score of 0.709±0.043 and a JAC (Jaccard Index) score of 0.551±0.053 when outlining the prostate. In the bladder adjacent to the prostate, the Dice score was 0.748±0.036 and the JAC score was 0.598±0.046. For the femoral head_L and femoral head_R, their Dice scores were both around 0.8, and JAC scores were around 0.7. The Dice score for the rectum was relatively low, at 0.634±0.045, and the JAC score was 0.466±0.051.
For the lungs, as shown in Figure 3 and Figure 5, SAM segment everything mode produced Dice scores around 0.86 for both lung_L and lung_R, and JAC scores around 0.75. The heart had a Dice score of 0.671±0.031 and a JAC score of 0.506±0.032. The spinal cord had a relatively low Dice score of 0.457±0.038, and a JAC score of 0.297±0.029. For the esophagus, SAM segment everything mode was unable to outline or recognize it, resulting in Dice and JAC scores of around 0.
For the gastrointestinal, as shown in Figure 4 and Figure 6, SAM segment everything mode produced a Dice score of 0.856±0.021 and a JAC score of 0.750±0.032 when outlining the liver. The results for kidney_L and kidney_R were similar, with Dice scores around 0.8 and JAC scores around 0.7. The spinal cord results were consistent with those of the lung, with Dice and JAC scores of about 0.45 and 0.3 respectively. The Dice score for the stomach was relatively low, at 0.319±0.027, and the JAC score was 0.190±0.021. For the large bowel and small bowel, SAM segment everything mode was unable to outline or recognize them, resulting in Dice and JAC scores of around 0.
For the head & neck, as shown in Figure 4 and Figure 6, SAM segment everything mode produced a Dice score of 0.896±0.018 and a JAC score of 0.812±0.030 when outlining the brain. The mandible had a Dice score of 0.868±0.022, and a JAC score of 0.768±0.033. The spinal cord had similar results to other sites, with Dice and JAC scores of about 0.4 and 0.3 respectively. For the parotid_L, parotid_R, cochlea_L, and cochlea_R, SAM segment everything mode was unable to outline or recognize them, resulting in Dice and JAC scores of around 0. All results were summarized in Table 1 (the third and fifth column).
@@ 4.3 Impact of box prompt
Overall, after introducing the manual box prompt, there was an improvement in Dice and JAC results for most organs. Some organs that could not be identified in SAM segment everything mode were now recognizable, although the resulting Dice and JAC scores were low. However, a few small organs in the head & neck remained unrecognizable even after the box prompt was employed.
@@
For the prostate, as shown in Figure 3 and Figure 5, the manual box prompt mode resulted in a Dice score of 0.883±0.030 and a JAC score of 0.792±0.049 when outlining the prostate. The bladder had a Dice score of 0.873±0.037, and a JAC score of 0.776±0.059. For femoral head_L and femoral head_R, their Dice scores each increased by around 0.1, and their JAC scores each increased by about 0.17. The rectum’s Dice score increased to 0.785±0.031, and the JAC score to 0.648±0.047.
@@
For the lungs, as depicted in Figure 3 and Figure 5, manual box prompt mode improved the Dice scores for both lung_L and lung_R by about 0.09, and the JAC scores by about 0.15. The spinal cord had a Dice score of 0.760±0.032, and a JAC score of 0.615±0.047-. The heart had an improvement with Dice scores and JAC scores increasing by about 0.16 and 0.21, respectively. The esophagus was recognizable under manual box prompt mode, but both Dice and JAC scores were low, at about 0.56 and 0.39, respectively.
@@
For the gastrointestinal, as shown in Figure 4 and Figure 6, the manual box prompt mode yielded a Dice score of 0.930±0.011 and a JAC score of 0.871±0.019 when outlining the liver. For kidney_L and kidney_R, the Dice and JAC scores both increased by about 0.09 and 0.13 respectively. The spinal cord saw an increase of around 0.3 for both the Dice and JAC scores. For the stomach, the Dice and JAC scores increased by about 0.26 and 0.21 respectively. The large bowel was recognizable under manual box prompt mode, but both Dice and JAC scores were low, at around 0.07 and 0.04 respectively. For the small bowel, manual box prompt mode recognized it, with both Dice and JAC values being around 0.15 and 0.09 respectively.
@@
For the head & neck, as shown in Figure 4 and Figure 6, the manual box prompt mode resulted in a Dice score of 0.938±0.012 and a JAC score of 0.883±0.021 when outlining the brain. The mandible had a Dice score of 0.915±0.021 and a JAC score of 0.843±0.039. The spinal cord had similar improvements as other regions, with increases of around 0.29 and 0.26 for the Dice and JAC scores respectively. For parotid_L and parotid_R, they were recognizable under manual box prompt mode, but both Dice and JAC scores were low, at around 0.57 and 0.40 respectively. For cochlea_L and cochlea_R, manual box prompt mode was still unable to recognize them, with both Dice and JAC values being around 0. All results were summarized in Table 1 (the fourth and sixth column).
@@
Figure 3: Boxplot (minimum, first quartile, median, third quartile, and maximum, respectively) of Dice coefficients of OARs between the ground clinical delineation and the SAM auto-segmented ones from three different experiments of the testing cases for prostate and lung. SAM segment everything and SAM manual box prompt correspond to 1 (grey) and 2 (red) in the figure.
@@
Figure 4: Boxplot (minimum, first quartile, median, third quartile, and maximum, respectively) of Dice coefficients of OARs between the ground clinical delineation and the SAM auto-segmented ones from three different experiments of the testing cases for gastrointestinal and head&neck. SAM segment everything and SAM manual box prompt correspond to 1 (grey) and 2 (red) in the figure.
@@
Figure 5: Boxplot (minimum, first quartile, median, third quartile, and maximum, respectively) of Jaccard coefficients of OARs between the ground clinical delineation and the SAM auto-segmented ones from three different experiments of the testing cases for prostate and lung. SAM segment everything and SAM manual box prompt correspond to 1 (grey) and 2 (red) in the figure.
@@
Figure 6: Boxplot (minimum, first quartile, median, third quartile, and maximum, respectively) of Jaccard coefficients of OARs between the ground clinical delineation and the SAM auto-segmented ones from three different experiments of the testing cases for gastrointestinal and head&neck. SAM segment everything and SAM manual box prompt correspond to 1 (grey) and 2 (red) in the figure.
Table 1: Dice and Jaccard scores of reference ROIs segmented by SAM in the prostate, lung, gastrointestinal(GI), and head&neck sites
Case Sites Reference ROIs Dice Jaccard
SAM-Everything SAM-Box SAM-Everything SAM-Box
Prostate Prostate 0.709±0.043 0.883±0.030 0.551±0.053 0.792±0.049
Bladder 0.748±0.036 0.873±0.037 0.598±0.046 0.776±0.059
Femoral head_L 0.807±0.028 0.908±0.023 0.678±0.040 0.832±0.038
Femoral head_R 0.799±0.038 0.910±0.029 0.669±0.056 0.836±0.048
Rectum 0.634±0.045 0.785±0.031 0.466±0.051 0.648±0.047
Lung Lung_L 0.862±0.025 0.943±0.016 0.758±0.039 0.892±0.028
Lung_R 0.860±0.027 0.951±0.015 0.756±0.041 0.906±0.027
Heart 0.671±0.031 0.839±0.023 0.506±0.032 0.723±0.035
Spinal cord 0.457±0.038 0.760±0.032 0.297±0.029 0.615±0.047
Esophagus 0 0.563±0.041 0 0.393±0.040
GI Liver 0.856±0.021 0.930±0.011 0.750±0.032 0.871±0.019
Kidney_L 0.813±0.029 0.904±0.023 0.687±0.039 0.826±0.037
Kidney_R 0.817±0.022 0.906±0.026 0.691±0.031 0.829±0.043
Spinal cord 0.441±0.039 0.733±0.020 0.283±0.027 0.579±0.028
Stomach 0.319±0.027 0.578±0.032 0.190±0.021 0.407±0.030
Small bowel 0 0.159±0.041 0 0.086±0.026
Large bowel 0 0.068±0.036 0 0.036±0.023
Head&Neck Brain 0.896±0.018 0.938±0.012 0.812±0.030 0.883±0.021
Mandible 0.868±0.022 0.915±0.021 0.768±0.033 0.843±0.039
Spinal cord 0.393±0.032 0.672±0.040 0.243±0.027 0.507±0.047
Parotid_L 0 0.573±0.026 0 0.401±0.023
Parotid_R 0 0.570±0.036 0 0.399±0.031
Cochlea_L 0 0 0 0
Cochlea_R 0 0 0 0
@@ 5 Discussion
In this section, we analyze the current performance of SAM segmentation in radiation oncology and expand on the possible future perspectives, practical applications, and prospective improvements of SAM in the realm of radiation oncology, based on the detailed results and insights from our study.
@@ 5.1 Current performance in radiation oncology-related segmentation
As for the Dice and Jaccard results, under the SAM "segment everything" mode, the auto segmentation outcomes were satisfactory for the prostate’s prostate, bladder, femoral head_L, and femoral head_R with (DICE: 0.7$\sim$0.8, JAC: 0.5$\sim$0.7), while they were less desirable for the rectum (DICE: $\sim$0.6, JAC: $\sim$0.4). For the lung, auto-segmentation for the lung_L and lung_R (DICE: 0.8$\sim$0.9, JAC: 0.7$\sim$0.8) were relatively better, yet less favorable for the heart (DICE: $\sim$0.6, JAC: $\sim$0.5) and spinal cord (DICE: $\sim$0.5, JAC: $\sim$0.3), with the esophagus not being recognized. The auto-segmentation outcomes for the gastrointestinal’s liver, kidney_L and kidney_R (DICE: 0.8$\sim$0.9, JAC: 0.7$\sim$0.8) were relatively better, whereas those for the stomach (DICE: $\sim$0.3, JAC: $\sim$0.2) were less satisfactory, with the large bowel and small bowel not recognized. For the head & neck, the brain and mandible (DICE: 0.8$\sim$0.9, JAC: 0.7$\sim$0.8)were better segmented, whereas the parotid_L, parotid_R, cochlea_L, and cochlea_R were not recognized. When comparing the results of segmentation across different sites, the "segment everything" mode in SAM performs better for the prostate and lung, but less satisfactory for the gastrointestinal and head & neck. If the volume and clarity of an organ are taken into account, it can be observed that the model performs better at delineating organs with clear boundaries and larger volumes, such as the liver and brain, and less satisfactory for organs with indistinct boundaries and smaller volumes, such as the parotid and cochlea, which is in general agreement with the experiences of manual delineation. The aforementioned results demonstrate that SAM’s performance in radiotherapy auto-delineation mirrors the clinical experience of human delineation, considering the variation across different sites and OARs. Importantly, SAM’s auto-segmentation is achieved by only one pre-trained model, suggesting that SAM exhibits robust generalizability for automatic delineation regarding different sites in radiotherapy.
@@
Upon the inclusion of the box prompt, SAM’s performance in auto-segmentation for radiotherapy showed further improvement. For most organs within the four reference sites, Dice and JAC scores have risen by 0.1 to 0.5, and previously unrecognized OARs such as the esophagus and parotid could be identified. This suggests that the box prompt, overall, is effective in enhancing SAM’s performance in radiotherapy segmentation, and future research could consider employing diverse prompt methods for further improvement. However, some OARs with vague boundaries, such as large bowel, small bowel and cochlea, were still not well recognized. This limitation is true for all auto-segmentation algorithms based on CTs. SAM is sensitive to the clarity of the OAR boundaries in radiotherapy delineation images; for some OARs with unclear boundaries, multi-modality images such as MRI could be considered to assist in the segmentation of some OARs. In addition, 3D segmentation is required for clinical radiotherapy, which warrants further research.
@@
Given the existing results, the SAM model demonstrates good generalization capabilities consistent with manual radiotherapy delineation, with a single model capable of executing all included test delineations for radiotherapy, and providing an effective prompt technology path for enhancing the performance of radiotherapy delineation. Judging from the Dice scores, SAM is able to meet the accuracy in the majority of OARs segmentation tasks required for clinical radiotherapy with Dice score higher than 0.7 when compared with manual delineation [92, 93]. Though parts of the OARs’ segmentation results from SAM are lower than 0.7, considering that the primary training of the SAM model largely relies on daily computer vision data, rather than professional clinical imaging data, the results are still acceptable, if not quite impressive [18]. This also suggests that future work can further fine-tune the SAM model in radiotherapy auto-segmentation through the incorporation of more clinical imaging data.